# Energy-Efficient Channel Decoding for Wireless Federated Learning: Convergence Analysis and Adaptive Design

### v1

@@ V-B Proof of Lemmas
$\mathbb{E}\|\mathbf{w}^{k}_{r,E}-Q^{\prime}(\mathbf{w}_{r})\|^{2}\leq\eta^{2}E(\sigma_{L}^{2}+\sum_{e=1}^{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}).$ (41)
Combining (40) and (41), we can conclude Lemma 4.
5) Proof of Lemma 5
Due to communication errors, the received DNN model for local training will contain random errors and differ from the global model, as shown in Fig. 11. In the following, we analyze the effect of the BER on the model error.
(a) BER=1E-2
(b) BER=1E-3
(c) BER=1E-4
Figure 12: Validation of the model error formula.
For a quantization scheme using $N$ bits to represent the weight $w$, the probability of having $x$ bit errors is given by
$P(X=x)=C_{N}^{x}b^{x}(1-b)^{N-x}.$ (42)
Since $b$ is much smaller than 1, the following relations can be established:
$P(X=0)\gg P(X=1)\gg P(X=2)\gg...\gg P(X=8).$ (43)
Therefore, the focus is on the case with at most one error bit.
The bit error can occur at various positions, ranging from the Least Significant Bit (LSB) to the Most Significant Bit (MSB), leading to different levels of error. For instance, for unsigned quantization, if the erroneous bit appears in the MSB and flips from 0 to 1, the error would be $+2^{N-1}\Delta$, where $\Delta=\frac{max(\mathbf{w})-min(\mathbf{w})}{2^{N}-1}=\frac{range(\mathbf{w})}{2^{N}-1}$, denoting the bin value for a given quantization. Alternatively, if the MSB flips from 1 to 0, the quantified error will be $-2^{N-1}\Delta$. The results for the case that the error occurs at the LSB will be $+2^{0}\Delta$ and $-2^{0}\Delta$. Additionally, it is assumed that the numbers of 0 and 1 in the bitstreams are similar. Based on this analysis, we can determine all the cases of bit flipping, and calculate the resulting values of the noisy weight with the corresponding probabilities, as shown in Table IV.
TABLE IV: Probability mass function of $(w)^{\prime}$.
$(w)^{\prime}_{i}$ $w+2^{0}\Delta$ $w-2^{0}\Delta$ $w+2^{1}\Delta$ $w-2^{1}\Delta$
$P_{i}$ b/2 b/2 b/2 b/2
$(w)^{\prime}_{i}$ $w+2^{N-1}\Delta$ $w-2^{N-1}\Delta$ $w$
[…]
Based on Table IV, the mean and variance of the received noisy weight $(w)^{\prime}$ can be calculated as follows:
$\mathbb{E}((w)^{\prime})=\sum_{i}((w)^{\prime}_{i}\cdot P_{i})=w,$ (44)
$\mathbb{E}||(w)^{\prime}-w||^{2}$ $=\sum_{i}[((w)^{\prime}_{i}-w)^{2}\cdot P_{i}]$
$=\frac{4^{N}-1}{3(2^{N}-1)^{2}}\cdot b\cdot range(\mathbf{w})^{2}.$ (45)
The above equations indicate that, for a weight parameter in the DNN model, the mean value of the noisy parameter is equal to the broadcasted perfect parameter. Furthermore, the variance between the two parameters is a function of the number of quantization bits, the BER level, and the range of the model parameters. For a $d-$dimensional model $\mathbf{w}$, i.e., $\mathbf{w}=[w_{1},w_{2},...,w_{d}]$, the following equations can be derived:
$\mathbb{E}||(\mathbf{w})^{\prime}-\mathbf{w}||^{2}$ $=\sum_{j=1}^{d}\mathbb{E}||(w_{j})^{\prime}-w_{j}||^{2}$
$=\frac{(4^{N}-1)d}{3(2^{N}-1)^{2}}\cdot b\cdot range(\mathbf{w})^{2}.$ (46)
To verify the theoretical results of Eq.(46), simulations were conducted using an 8-bit vanilla-CNN on the Fashion-MNIST dataset. The experimental settings are described in Section IV-A. Three different BER conditions were simulated, and the model errors obtained from the simulation are plotted in Fig. 12. The expectation on the left-hand side (LHS) of Eq.(46) corresponds to different cases of errors. The model errors predicted using the proposed formula on the RHS of Eq.(46) are also plotted in the same figure. It can be seen that, for different BER conditions and different communication rounds, the results obtained from the theoretical formula closely match those obtained from simulation, confirming the accuracy of the derived Eq.(46).
If we denote $M$ as the maximum absolute value of all weight elements, i.e. $,M=max(|w_{1}|,|w_{2}|,$$,|w_{d}|)$, we have $range(\mathbf{w})\leq 2M$. The upper bound of Eq.(46) can be obtained as follows:
@@
$\mathbb{E}||(\mathbf{w})^{\prime}-\mathbf{w}||^{2}\leq\frac{4d(4^{N}-1)}{3(2^{N}-1)^{2}}\cdot b\cdot M^{2},$ (47)
which concludes the proof of Lemma 5.

### v2

@@ V-B Proof of Lemma 2
$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L(L+2)}{2}\mathbb{E}\|\mathbf{w}_{r}-\mathbf{\tilde{w}}_{r}\|^{2}\\ &+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &+\frac{L(L+1)E\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (34)
Summing (34) over $r=0,\cdots,R-1$, rearranging the terms, and multiplying $\frac{2}{\eta RE}$ on both sides, we obtain the following expression:
$\begin{split}&\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-\mathbb{E}f(\mathbf{w}_{R}))}{\eta RE}\\ &+\frac{L(L+2)}{\eta RE}\sum_{r=0}^{R-1}\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}+\frac{1}{\eta RE}\sum_{r=0}^{R-1}\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}\\ &+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}.\end{split}$ (35)
Based on (23) and (21) respectively, we have
@@
$\begin{split}\mathbb{E}\left[||\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}||^{2}\right]&=\mathbb{E}\left[\mathbb{E}\left[||\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}||^{2}|\mathbf{w}_{r}\right]\right]\\ &=\frac{d(4^{N}-1)}{3(2^{N}-1)^{2}}\cdot b_{r}\left(1-b_{r}\right)^{N-1}\cdot M_{r}^{2},\end{split}$ (36)
$\begin{split}\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}&=\sum_{j=1}^{d}\mathbb{E}\left[|\delta(w_{r,d})|^{2}|{w}_{r,d}\right]\\ &\leq\sum_{j=1}^{d}\mathbb{E}\left[(b_{r}(1-b_{r})^{N-1}M)^{2}|{w}_{r,d}\right]\\ &\leq d(b_{r}(1-b_{r})^{N-1})^{2}M_{r}^{2}.\end{split}$ (37)
By substituting (36) and (37) into (35), and denote $f^{*}$ as the minimum global training loss, we have
$\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-f^{*})}{\eta RE}$
$+\frac{L(L+2)d(4^{N}-1)}{3(2^{N}-1)^{2}\eta RE}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}M_{r}^{2}$
$+\frac{d}{\eta RE}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}M_{r}^{2}$
$+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}.$ (38)
Denoting $\hat{M}=max(M_{0},M_{1},\cdots,M_{R-1})$, we have
$\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-f^{*})}{\eta RE}$
$+\frac{L(L+2)d\hat{M}^{2}(4^{N}-1)}{3(2^{N}-1)^{2}\eta RE}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}$
$+\frac{d\hat{M}^{2}}{\eta RE}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}$
$+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}.$ (39)